OpenStudy (sunshinehaggard):

What is the first step in solving ln(x − 1) = ln6 − lnx for x?

OpenStudy (isaidavila):

First remember soñé properties Ln(a/b)=lna-lnb

OpenStudy (isaidavila):

Ln(x-1)=Ln(6/x)

OpenStudy (isaidavila):

Since Ln are equal the numbers are also equal x-1=6/x

OpenStudy (isaidavila):

Now solve for x